OpenStudy (anonymous):

Find the equation of a line that is perpendicular to the line y=1/8x +8 and contains the point (-2,0).

OpenStudy (anonymous):

Perpendicular lines have their slopes as a flipped fraction of the original slope and opposite signed. A perpendicular line to an equation with a slope of (2/5) would have a slope of (-5/2)

OpenStudy (anonymous):

What is the slope of the equation you posted? Flip the fraction and make it negative to find the slope of the perpendicular equation.

OpenStudy (anonymous):

*the slope is always the number next to x

OpenStudy (anonymous):

The slope of the equation is (1/8) so a perpendicular slope would be -8

OpenStudy (anonymous):

-8x +8

OpenStudy (anonymous):

Ignore the +8, that's the y=intercept of the original equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now we use point-intercept form. y-(y1)=m(x-x1) Using the point (-2,0). 0 is y1, m is perpendicular slope (-8), and x is -2

OpenStudy (anonymous):

y-0=-8(x-(-2) solve the equation for y

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

trying to solve it.... a bit slow but getting there

OpenStudy (anonymous):

y=-8x +16

OpenStudy (anonymous):

y-0=-8(x-(-2)) y-0=-8(x+2) x-(-2) = x+2

OpenStudy (anonymous):

You are close, it should be y=-8x-16 (double checking)

OpenStudy (anonymous):

y=-8(x+2) y=-8x-16

OpenStudy (anonymous):

sure negative 8 times positive 2 is negative 16. got it. thx

OpenStudy (anonymous):

i always miss that

OpenStudy (anonymous):

People get confused with x-negative number, Minus a negative is a positive.

OpenStudy (anonymous):

Will remember that. thx.

OpenStudy (anonymous):

Thanks larrybox for simplifying this for me. you sure made it easy to understand.

OpenStudy (anonymous):

In the eqation y=5x+5, what would be the perpendicular slope?

OpenStudy (anonymous):

so what happens to the y intercept

OpenStudy (anonymous):

meant -5x

OpenStudy (anonymous):

Be careful, in this equation, the fraction is (5/1)

OpenStudy (anonymous):

So the perpendicular slope would be (-1/5)

OpenStudy (anonymous):

Nothing happens to the y-intercept, it just is not needed in calculation.

OpenStudy (anonymous):

you know what? I thought about that but changed my mind. silly me!

OpenStudy (anonymous):

Anything over 1 is itself. (5/1) 5, (-1/1) = -1

OpenStudy (anonymous):

maths maths maths! not my strongest point at all

OpenStudy (anonymous):

if the perpendicular slope for 1/8x = -8, then 5x = 5/1 would be -1/5

OpenStudy (anonymous):

i should start saying math is interesting and gradually will get there! Thx lbp